An investor invests 40 percent of his wealth in a risky asset with an expected rate of return of 0.15 and a variance of 0.04 and 60 percent in a T-bill that pays 6 percent. His portfolio's expected return and standard deviation are __________ and __________, respectively.

A.

0.096; 0.12

B.

0.295; 0.12

C.

0.795; 0.14

D.

0.114; 0.12

E.

0.096; 0.08
Answer #1

Given about risky assets,

expected rate of return E(r) = 0.15

Variance = 0.04

standard deviation SD(r) = Sqrt(variance) = 0.2

Investment weight wr = 40%

Risk free rate Rf = 0.06

investment is risk free T-Bill wf= 60%

So, expected return of the portfolio is weighted average return of its assets

E(p) = wr*E(r) + wf*Rf = 0.4*0.15 + 0.6*0.06 = 0.096

standard deviation of portfolio = wr*SD(r) = 0.4*0.2 = 0.08

So, option E is correct.